NPP Projection Closee 3/24 + 3/25 Admin



To do this evening/ was completed: Re-confirmed all of Tuesday's OCAP screenings are good to go.

Scheduled the upcoming week in LMS with a few exceptions. Started the ingest for Minecraft Movie and Eephus. Those playlists yet to be built.

REC played well last night.

NSF Playlist tech went well. We caught some kinks and those have mostly been corrected. Kat sent a recap email highlighting what we caught but I'll reiterate here:

The NSF24_THANKYOU slide needs to appear before the program holding slide and after the MAKETHEROAD slide that comes at the end of the program. I've updated the playlists to include this.

We found that the durations of THANKYOU slide and the MAKETHEROAD slide need to be changed. Thank You at 10 seconds and Road at 30, so I've made new DCPs of each. The Thank You slide I made by downloading the previous version's mxf file and trimming the content by 20 seconds in DCPoMatic. The MAKETHEROAD slide we made by screenshotting the slide in VLC and making that image 30 seconds long (couldn't simply increase mxf file length in DCPoMatic). The screenshot had a ratio of 1.79, I changed it to 1.85. I don't think this will have an effect on the QR code but noting it here. Also: these two new slides do not fully display their titles in their file names but you can distinguish them with their creation date "20250325". MAKETHEROAD slide is titled SlideNSF25V2Ma_XSN... All NSF playlists updated. Take a look.

The 1.85 lens/mask pack was failing to trigger 1.85 masking. When I looked at the pack I found the FILM masking cue in there instead of the regular masking cue so I've updated the pack.

We still need to make a new lens file for 1.90.

Tech'd Conquest with McKegg and Cristina. Mono 8.0 is the final decision. Ran loops afterward and cleaned the projectors.

 Presentation Issues:

Preshow Clips/Playlist:

Trailer Updates:

Special Events/ AV set up:

Ordering/Receiving (posters & trailers):

Ordering/Receiving (content): Keys for Flow, Novocaine, Death of a Unicorn, and Opus have been uploaded

Outgoing/Shipping (content):

Print/video Inspection:

Asset tracking doc updated/checked?:

Maintenance/cleaning:

Cue sheets updated/checked (every shift): Next week's cue sheets filled out with the exception of Eephus and The Friend. Formulas for those still need to be made. 

Marquee updated?:

Posters updated?:

Cleaning lights on for the night?:

Projection cart by entrance?:

Additional Notes:

Comments

Popular posts from this blog

NPP Projection Open 12/28

Projection Admin (and Close) 10/19

NPP Projection Report Friday 1/31 CLOSE Chris